# AlignXIE: Improving Multilingual Information Extraction by Cross-Lingual Alignment

### v1

@@ 3.1.2 IE Parallel Data Construction Pipeline
The latter two stages address the challenge of missing spans through rephrasing. This stage focuses on rephrasing the translated span. Consequently, if $i^{tgt}_{m}\notin s^{tgt}$, we prompt the LLM to find the span $i^{cor}_{m}$ in $s^{tgt}$ that semantically matches with the $i^{src}_{m}$. If $i^{cor}_{m}\in s^{tgt}$, $i^{tgt}_{m}$ will be rephrased as $i^{cor}_{m}$.
Stage 3: Sentence Rephrase
This stage focuses on the rephrasing of the translated sentence. If not all $i^{tgt}_{m}\in s^{tgt}$, we prompt the LLM to rephrase the $s^{tgt}$ based on the $s^{src}$ to ensure all $i^{tgt}_{m}$ are included.
Following Parekh et al. (2024), we evaluated the pipeline on the label projection benchmark WikiANN Pan et al. (2017), a multilingual IE dataset, and achieved average $99\%$ faithfulness in $10$ English-centric bilingual settings. The detailed results and settings are shown in Appendix A.1.
3.1.3 ParallelNER
@@
[…]
3.2 Multilingual IE Training Phase
Following the previous phase, we conducted multilingual IE instruction tuning with $46$ IE datasets in English and Chinese. The detailed dataset statistics are shown in Appendix B. We follow the instruction and completion format introduced in Section 2.
4 Experimental Settings
AlignXIE enhances IE cross-lingual transfer by aligning the schema and extraction process. Our extensive evaluation of IE cross-lingual transfer can be divided into two main dimensions: external-language transfer (cross-lingual IE), and internal-language transfer (multilingual IE).
4.1 Metrics
@@
Table 2: Results of cross-lingual evaluation in Multiconer22. Avgcross represents the average performance across $9$ unseen languages. indicates that the score of ChatGPT is reported by Lai et al., 2023. Supervised denotes the supervised method which is a strong baseline since the above four methods are under zero-shot setting.
@@ 4.2 Implementation Details
AlignXIE is fine-tuned based on Baichuan2-7B-Base Baichuan (2023). We utilize the LLaMA-Factory Zheng et al. (2024) for instruction tuning. We use LoRA Hu et al. (2022) for parameter-efficient fine-tuning and set the LoRA rank to $32$. The warmup ratio is set to $0.01$. The learning rates for these two phases are set to $3\times 10^{-4}$. We limit the sequence length to $4096$ and set the batch size to $256$. During the inference phase, we set the temperature to 00. All experiments are conducted on 8 x NVIDIA-A800 80G.
Dataset BERT YAYI-UIE IEPILE B2NER AlignXIE
@@
Table 3: Results of supervised evaluation on NER. Results of BERT on English benchmarks are from Wang et al. (2023a), and on Chinese benchmarks are from Li et al. (2020a).
4.3 Datasets
We followed the IEPILE Gui et al. (2024), YAYI-UIE Xiao et al. (2023), and B2NER Yang et al. (2024) to conduct a comprehensive evaluation . Overall, for English IE, we evaluate performance across $28$ benchmarks for NER, $10$ benchmarks for RE, $6$ benchmarks for ED, and $3$ benchmarks for EAE. For Chinese IE, we evaluate performance on $4$ benchmarks for NER, $5$ benchmarks for RE, $4$ benchmarks for ED, and $2$ benchmarks for EAE. For cross-lingual IE, we evaluate performance on the multilingual NER benchmark Multiconer22 Malmasi et al. (2022b).
We conduct supervised evaluations on various benchmarks: $23$ NER, $8$ RE, $3$ ED, and $3$ EAE benchmarks in English, as well as $2$ NER, $2$ RE, $2$ ED, and $2$ EAE benchmarks in Chinese. The detailed statistics and usage are presented in Appendix B. For the zero-shot evaluation, among NER benchmarks, we use $5$ English benchmarks from CrossNER Liu et al. (2021) and $2$ Chinese benchmarks, Weibo Peng and Dredze (2015) and Boson [^1] https://github.com/InsaneLife/. Among RE benchmarks, we adopt $2$ English benchmarks FewRel (Han et al., 2018), Wiki-ZSL (Chen and Li, 2021) and $3$ Chinese benchmarks, COAE2016222https://github.com/Sewens/COAE2016, IPRE (Wang et al., 2019), and SKE2020333https://aistudio.baidu.com/datasetdetail/177191. Among ED benchmarks, we use $3$ English benchmarks CrudeOil News Lee et al. (2021), RAMS Ebner et al. (2020), WikiEvents Li et al. (2021) and $2$ Chinese benchmarks FewFC (Zhou et al., 2021) and CCF law [^4] https://aistudio.baidu.com/projectdetail/4201483.
4.4 Baselines
We mainly use two types of baselines for a comprehensive comparison: Following InstructUIE Wang et al. (2023a), we first compare AlignXIE with BERT-Base Devlin (2018) methods. Additionally, we compare AlignXIE with the SoTA multilingual IE systems based on LLMs, including IEPILE Gui et al. (2024), YAYI-UIE Xiao et al. (2023), and B2NER Yang et al. (2024), which are based on natural language prompts. The three baselines exhibit varying degrees of unfairness when compared to AlignXIE . IEPILE and YAYI-UIE are fine-tuned based on the Baichuan2-13B, while B2NER uses Intern2LM-7B Cai et al. (2024).
For the cross-lingual evaluation, we compare with ChatGPT (evaluated by Lai et al., 2023), GLiNER Zaratiana et al. (2024), and B2NER. Furthermore, we further compare with the supervised baseline in Multiconer22 Malmasi et al. (2022b).
5 Results
5.1 Cross-Lingual Evaluation
To evaluate whether AlignXIE improves IE cross-lingual alignment, we conducted the cross-lingual evaluation on $9$ unseen languages of Multiconer22. We have also supplemented the evaluation in Chinese and English. Table 2 shows results. AlignXIE significantly outperforms ChatGPT and GLiNER in the cross-lingual setting, achieving a $30.17\%$ and $20.03\%$ improvement, respectively. For $4$ unseen languages used in B2NER that constitute more than $0.1\%$ of the general LLM pretraining corpus Touvron et al. (2023c), AlignXIE outperforms B2NER by $19.36\%$, with results comparable to B2NER in both Chinese and English, demonstrating a significant improvement in cross-lingual transfer despite lacking supervised training data as described in §4.4.
Dataset YAYI-UIE IEPILE AlignXIE
@@ 5.1 Cross-Lingual Evaluation
Surprisingly, despite not having any training data in $9$ unseen languages, AlignXIE has achieved performance comparable to the supervised baseline in Spanish and Turkish. The results demonstrate that AlignXIE exhibits excellent cross-lingual generalization for IE, which can be attributed to the enhancement of IE cross-lingual alignment.
Task Dataset BERT YAYI-UIE IEPILE AlignXIE
@@
Table 5: Results of supervised evaluation on EE. Results of BERT are from Xiao et al. (2023).
5.2 Supervised Evaluation
The results for NER, RE, EE (including ED and EAE) tasks are shown in Tables 3, 4, and 5 respectively. AlignXIE outperforms the multilingual SoTA baselines on most benchmarks for four tasks and ranks within the top-$2$ results across RE and ED benchmarks. In the English IE, AlignXIE has achieved significant average improvements of 3.03 and 2.92 F1 points on the RE and ED tasks, with an improvement of $7.85$ points on the kbp37 of RE. In the Chinese IE, AlignXIE has consistently achieved SoTA, with $4.12$ points average improvement over the SoTA baseline on the EAE task. Moreover, it can be observed that the model not only achieves significant improvements in NER but also exhibits even greater enhancements in RE, ED, and EAE. This indicates that the alignment phase enhances the multilingual NER capability through cross-lingual alignment, which improves the performance of all tasks by leveraging the foundational spotting capabilities of NER.
@@ 5.2 Supervised Evaluation
To further demonstrate that AlignXIE improves multilingual IE across all languages through IE cross-lingual alignment, we also compared it comprehensively with SoTA monolingual (English) IE systems, which are fine-tuning on LLMs including InstructUIE Wang et al. (2023a), UniversalNER Zhou et al. (2024), GoLLIE Sainz et al. (2024), KnowCoder Li et al. (2024), GLiNER Zaratiana et al. (2024), and GNER Ding et al. (2024). Appendix A.2 shows a detailed comparison with all baselines. Compared to $10$ SoTA methods in the supervised setting, it achieves top-$2$ results on $14$ (of $23$) English NER benchmarks, indicating substantial enhancements. Specifically, when compared to other code-based IE systems, such as KnowCoder and GoLLIE, our model still achieves a notable improvement. This observation underscores the effectiveness of cross-lingual alignment of AlignXIE . This indicates that our cross-lingual alignment not only avoids conflicts between different languages but also enhances overall multilingual IE performance by leveraging cross-language alignment.
Task Dataset GPT-4 YAYI-UIE IEPILE AlignXIE

### v2

@@ 4.2 IE Parallel Data Construction Pipeline
Stage 2: Span Rephrase   Stage 2 and 3 tackle the missing spans issue by rephrasing. In Stage 2, $i^{tgt}_{k}$ corresponding to $i^{src}_{k}$ is identified from $s^{tgt}$, similar to the CLaP Parekh et al. (2024), which allows the rephrasing of the previously translated ones to further address the missing spans issue that persists after Stage 1. Specifically, if $i^{tgt}_{m}\notin s^{tgt}$, we prompt the LLM to find the span $i^{tgt-cor}_{m}$ in $s^{tgt}$ that semantically matches with $i^{src}_{m}$. If $i^{tgt-cor}_{m}\in s^{tgt}$, $i^{tgt}_{m}$ will be rephrased as $i^{tgt-cor}_{m}$.
Stage 3: Sentence Rephrase   However, due to the inherent issue of error accumulation in the traditional sentence-then-span approach employed by CLaP-like methods, we introduce Stage 3. In this stage, we rephrase the translated sentence that is a span-then-sentence strategy, thereby mitigating the error propagation of the sentence-then-span approach and addressing the missing spans issue that remains unresolved. Specifically, if not all $i^{tgt}_{m}\in s^{tgt}$, we prompt the LLM to rephrase $s^{tgt}$ based on $s^{src}$ to ensure all $i^{tgt}_{m}$ are included.
This pipeline is capable of generating IE parallel data across any two languages, as demonstrated in Algorithm 1. By integrating three translation strategies, parallel processing, sentence-then-span, and span-then-sentence, we have achieved unprecedented performance on the label projection task, approaching an accuracy rate of nearly 100%. Following Parekh et al. (2024), we evaluate the pipeline on the label projection benchmark WikiANN Pan et al. (2017), a multilingual IE dataset, and achieve an average $99\%$ on the faithfulness evaluation in $10$ English-centric bilingual settings. Appendix B shows the detailed settings and results.
4.3 ParallelNER
@@
Table 2: Results of cross-lingual evaluation in Multiconer22. Avgcross represents the average performance across $9$ unseen languages. ${}^{\textbf{\textdagger}}$ indicates that the score is reported by Lai et al. (2023). ${}^{\textbf{\textdaggerdbl}}$ indicates that the score is from our implementation. indicates that the result of GLiNER in Chinese is under the cross-lingual setting. Supervised denotes the supervised method, which is a strong baseline since the above methods are under the zero-shot setting.
5 Experimental Settings
@@ 5.1 Implementation Details
KnowCoder-X is fine-tuned based on Baichuan2-7B-Base Baichuan (2023) using LLaMA-Factory Zheng et al. (2024). We apply LoRA Hu et al. (2022) with a LoRA rank of $32$ for parameter-efficient fine-tuning. The warmup ratio is set to $0.01$, and the learning rate for both phases is $3\times 10^{-4}$. The sequence length is limited to $4096$, and the batch size is $256$. During inference, the temperature is set to 00. All experiments are conducted on 8 x NVIDIA A8000 80G GPUs.
5.2 Datasets
We follow the IEPILE Gui et al. (2024), YAYI-UIE Xiao et al. (2023), and B2NER Yang et al. (2024) to conduct a comprehensive evaluation. Overall, for English IE, we evaluate performance across $28$ benchmarks for NER, $10$ benchmarks for RE, $6$ benchmarks for ED, and $3$ benchmarks for EAE. For Chinese IE, we evaluate performance on $4$ benchmarks for NER, $5$ benchmarks for RE, $4$ benchmarks for ED, and $2$ benchmarks for EAE.
Specifically, for cross-lingual evaluation, we evaluate performance on the multilingual NER benchmark Multiconer22 Malmasi et al. (2022b) and the low-resource African language NER benchmark MasakhaNER 2.0 Adelani et al. (2022). Moreover, we conduct the supervised evaluation on various benchmarks: $23$ NER, $8$ RE, $3$ ED, and $3$ EAE benchmarks in English, as well as $2$ NER, $2$ RE, $2$ ED, and $2$ EAE benchmarks in Chinese. Appendix G presents the detailed statistics and usage.
We only train on the supervised evaluation IE datasets of the three main baselines to ensure a fair comparison, without additional training data.
5.3 Baselines
To enable a fair and comprehensive comparison, we mainly use two types of baselines: Following B2NER Yang et al. (2024), we first compare KnowCoder-X with BERT-Base Devlin (2018). Additionally, we compare KnowCoder-X with the SoTA systems for English and Chinese IE based on LLMs, including IEPILE Gui et al. (2024), YAYI-UIE Xiao et al. (2023), and B2NER Yang et al. (2024), which are our main baselines and based on natural language prompts. The three baselines exhibit varying degrees of unfairness compared to KnowCoder-X. IEPILE and YAYI-UIE are fine-tuned on the larger backbone Baichuan2-13B, while B2NER utilizes the more advanced backbone Intern2LM-7B  Cai et al. (2024).
For the cross-lingual evaluation, we further compare KnowCoder-X with ChatGPT (from Lai et al., 2023) and GLiNER Zaratiana et al. (2024).
6 Results
Our extensive evaluation of IE cross-lingual transfer can be divided into two main dimensions: the cross-lingual evaluation for external-language transfer in unseen languages (§6.1), and the supervised evaluation for internal-language transfer between English and Chinese (§6.2). We additionally conduct the zero-shot evaluation for internal language transfer. Appendix D shows the results.
6.1 Cross-Lingual Evaluation
To evaluate whether KnowCoder-X improves IE cross-lingual alignment, we conduct the cross-lingual evaluation on $9$ unseen languages of Multiconer22. We have also supplemented the evaluation in Chinese and English. Table 2 shows results. KnowCoder-X significantly outperforms ChatGPT and SoTA in the cross-lingual setting, achieving a $30.17\%$ and $20.03\%$ improvement, respectively. For $4$ unseen languages used in B2NER that constitute more than $0.1\%$ of the general LLM pretraining corpus Touvron et al. (2023c), KnowCoder-X outperforms B2NER by $19.36\%$, with results comparable to B2NER in both Chinese and English, demonstrating a significant improvement in cross-lingual transfer. Surprisingly, despite no training data in $9$ unseen languages, KnowCoder-X has achieved performance comparable to the supervised method in Spanish and Turkish that significantly outperforms the SoTA.
Appendix C presents the evaluation results on MasakhaNER 2.0 with $20$ African low-resource languages. The results demonstrate that KnowCoder-X exhibits strong cross-lingual performance and can extend to a wide range of languages, particularly those with limited resources.
Dataset BERT YAYI-UIE IEPILE B2NER KnowCoder-X
@@ 6.1 Cross-Lingual Evaluation
Table 3: Results of supervised evaluation on NER. Results on Chinese benchmarks are from Li et al. (2020a).
6.2 Supervised Evaluation
The results for NER, RE, EE (including ED and EAE) tasks are shown in Tables 3, 4, and 5 respectively. KnowCoder-X outperforms the multilingual SoTA baselines on most benchmarks for four tasks and ranks within the top-$2$ results across all RE and ED benchmarks. In the English IE, KnowCoder-X has achieved significant average improvements of $3.03$ and $2.92$ F1 points on the RE and ED tasks, with an improvement of $7.85$ points on the kbp37 of RE. In the Chinese IE, KnowCoder-X has consistently achieved SoTA, with $4.12$ points average improvement over the SoTA baseline on the EAE task. Moreover, it can be observed that the model not only achieves significant improvements in NER but also exhibits even greater enhancements in RE, ED, and EAE. This indicates that the alignment phase enhances the multilingual NER capability through cross-lingual alignment, which improves the performance of all tasks by leveraging the foundational spotting capabilities of NER.
Dataset YAYI-UIE IEPILE KnowCoder-X
@@ 6.2 Supervised Evaluation
Task Dataset BERT YAYI-UIE IEPILE KnowCoder-X
@@
Table 5: Results of supervised evaluation on EE.
@@
To further demonstrate that KnowCoder-X improves multilingual IE across all languages through IE cross-lingual alignment, we also compared it comprehensively with SoTA monolingual (English) IE systems, which are fine-tuned on LLMs including InstructUIE Wang et al. (2023), UniversalNER Zhou et al. (2024), GoLLIE Sainz et al. (2024), KnowCoder Li et al. (2024b), GLiNER Zaratiana et al. (2024), and GNER Ding et al. (2024). Appendix E shows a detailed comparison with all baselines. Compared to $10$ SoTA methods in the supervised setting, it achieves top-$2$ results on $14$ (of $23$) English NER benchmarks, indicating substantial enhancements. Specifically, when compared to other code-based IE systems, such as KnowCoder and GoLLIE, KnowCoder-X still achieves a notable improvement. This observation underscores the effectiveness of cross-lingual alignment of KnowCoder-X. This indicates that our cross-lingual alignment not only avoids conflicts between different languages but also enhances overall multilingual IE performance by leveraging cross-lingual alignment. Appendix F shows a comparison of the computational resources.
Dataset Text-Based Code-Based
ACE 2005 Walker et al. (2006) 84.93 86.07
AnatEM Pyysalo and Ananiadou (2014) 89.63 90.14
BC2GM Kocaman and Talby (2020) 82.53 83.17
BC5CDR Li et al. (2016) 90.03 90.17
CoNLL 2003 Sang and Meulder (2003) 94.67 94.93
WNUT 2017 Derczynski et al. (2017) 66.14 66.03
ResumeNER Zhang and Yang (2018) 94.31 95.37
MSRA Levow (2006) 92.97 94.03
Avg 86.90 87.49
Table 6: Ablation study on code-based multilingual IE.
Model NER RE ED EAE Avg
KnowCoder-X (w/o align) 85.06 68.67 74.12 74.00 75.46
KnowCoder-X 85.63 70.38 75.37 74.87 76.56
Table 7: Ablation study on IE alignment training.
6.3 Ablation Study
We conduct comprehensive ablation experiments to investigate whether the code-based multilingual IE and IE cross-lingual alignment phase of KnowCoder-X contribute to the performance.
Code-Based Multilingual IE   Table 6 shows results. We evaluate the text-based Wang et al. (2023) and code-based multilingual IE method in the supervised setting. We use $6$ English NER datasets, ACE 2005, AnatEM, BC2GM, BC5CDR, CoNLL 2003, WNUT 2017, and $2$ Chinese NER datasets, MSRA, ResumeNER to conduct instruction tuning on the Baichuan2-7B for supervised evaluation. We remove class comments in our schemas to ensure a fair comparison. The code-based method outperforms by $0.59$ points, with improvements exceeding $1.00$ points in $3$ datasets, particularly in the Chinese datasets MSRA and ResumeNER. The results show that the code can mitigate schema differences between languages, thereby enhancing the cross-lingual alignment to improve the performance of multilingual IE.
IE Alignment Training   To show whether the IE cross-lingual alignment phase contributes to the performance, we further conduct an ablation study by removing the cross-lingual alignment phase of KnowCoder-X and denoting it as KnowCoder-X (w/o align phase). Table 7 shows the average results in the supervised setting on $4$ tasks across $45$ benchmarks. Compared to KnowCoder-X (w/o align phase), the results indicate that KnowCoder-X achieves a substantial average improvement of $1.1$ points across four tasks. This suggests that the IE cross-lingual alignment phase enhances the overall performance of multilingual IE, demonstrating its contribution to facilitating cross-lingual transfer within the multilingual IE framework.
Language Text-Based Code-Based Code-Based (w/ com.)
German 35.16 38.69 48.02
Spanish 43.19 48.93 52.16
Dutch 39.16 42.19 46.79
Russian 33.12 36.19 39.68
Bengali 18.23 17.59 23.95
Persian 23.45 24.69 28.54
Hindi 21.71 23.91 28.13
Korean 28.31 28.65 31.42
Turkish 32.36 35.16 41.16
Avg 30.52 32.89 37.76
Table 8: Analysis on cross-lingual transfer.
6.4 Analysis on Cross-Lingual Transfer
To further investigate the impact of code-based unified schema representation in cross-lingual IE, we conduct cross-lingual evaluation finetuned on all NER datasets under three settings: text-based, code-based, and code-based (with comments). Table 8 shows the results. It indicates that adopting the code-based multilingual IE method can enhance cross-lingual performance by $2.37\%$. This improvement can be attributed to the standardized representation of the same ontology across different languages provided by the code-based multilingual IE method. Additionally, the integration of guideline information further improves cross-lingual performance by $4.85\%$, particularly in low-resource languages such as Bengali and Persian. The code-based multilingual IE effectively facilitates the integration of guideline information, significantly enhancing cross-lingual IE performance.
7 Conclusion
We introduce KnowCoder-X, a code-based multilingual IE model that significantly enhances cross-lingual transfer through unified schema representation and IE cross-lingual alignment training. KnowCoder-X utilizes Python classes to represent schemas uniformly across languages, ensuring semantic consistency in multilingual IE. Additionally, the IE cross-lingual alignment phase leverages high-quality IE parallel datasets ParallelNER, which are generated by our proposed LLM-based pipeline, to boost multilingual generalization. Comprehensive experiments demonstrate that KnowCoder-X achieves SoTA in Chinese and English IE under various settings, with remarkable cross-lingual capability.
Limitation
Expanding training to include languages beyond English and Chinese remains an area for future exploration. At the same time, incorporating additional languages would provide valuable insights into the cross-lingual alignment mechanisms within multilingual IE systems. Additionally, addressing domain specificity will be a key focus of future research. Specialized vocabulary and domain-specific language use present significant challenges for cross-lingual transfer, making it essential to overcome these obstacles in upcoming work. Introducing parallel data from other IE tasks during the cross-lingual alignment phase also offers promising avenues for further investigation.
Appendix A Description Generation
We provide a detailed explanation of the description generation process using the NER task as an example. The generation process consists of two main phases, i.e., Description Initialization and Description Polish:
  • Description Initialization: In this phase, we sample some entities to instruct the LLMs to summarize an initial description of these entities. The details are as follows: 1) Randomly sample $10$ entities from the training set. 2) Format the $10$ entities into the description initialization prompt template as shown in Tables 20, instructing the LLM to summarize and generate an initial description.
  • Description Polish: Based on the description generated by the first phase, this phase further instructs the LLM to polish the description based on a set of entities iteratively. The details are as follows: 1) Randomly sample $20$ entities from the training set for iterative refinement. 2) Iterate over each entity from the entity set above: Format the current entity with the current description into the description polish template as shown in Table 21, instructing the LLM to evaluate whether the current description adequately covers the example semantically: If the current description adequately covers the entity, the LLM will not change the description. If the current description fails to encompass the entity, the LLM will revise the inappropriate parts of the description, ensuring the description covers the current entity.
Description Initialization often struggles to accurately represent and capture the characteristics of each entity. To address this, we introduce Description Polish, a method designed to refine the details of the description, thereby enabling it to summarize the entities within a given entity type more effectively.
Appendix B Pipeline Evaluation
We evaluate our pipeline on the label projection benchmark WikiANN Pan et al. (2017), which is a multilingual NER dataset. We use $10$ languages, including Bengali (bn), German (de), Spanish (es), Persian (fa), Hindi (hi), Korean (ko), Dutch (nl), Russian (ru), Turkish (tr), and Chinese (zh), and randomly sample $1,000$ samples for each language. We compare our pipeline with CLaP Parekh et al. (2024), Awesome-Align Dou and Neubig (2021), and EasyProject Chen et al. (2023). We mainly evaluate faithfulness following the setting of CLaP, measured as the percentage of instances where translated labels appear in the translated sentence, as accuracy evaluation requires costly native speaker rankings across methods, incurring high costs.
Language Awesome-Align EasyProject CLaP KnowCoder-X
bn 92 98 93 99
de 99 97 79 99
es 99 99 84 100
fa 96 99 72 100
hi 93 36 90 99
ko 96 93 64 99
nl 99 100 85 100
ru 97 99 66 93
tr 98 98 94 99
zh 92 92 60 99
Avg 96 91 79 99
Table 9: Results of label projection evaluation.
Appendix C Cross-Lingual Evaluation on Low Resource Languages
Table 10 shows results on the MasakhaNER 2.0. KnowCoder-X achieves the best performance across a wider variety of low-resource languages. Compared to IEPILE and GLiNER, KnowCoder-X achieves improvements of $11.41\%$ and $27.11\%$, respectively. These findings highlight the remarkable cross-lingual performance of KnowCoder-X on low-resource languages. The results demonstrate that KnowCoder-X exhibits excellent cross-lingual generalization for IE, which can be attributed to the enhancement of IE cross-lingual alignment.
Language YAYI-UIE IEPILE GLiNER KnowCoder-X
Bamanankan 31.19 29.00 28.85 32.97
Ghomala 32.03 32.36 34.71 32.97
Éwé 63.01 57.83 64.64 66.05
Fon 30.05 29.73 30.58 31.86
Hausa 47.68 43.06 42.04 48.35
Igbo 37.14 37.54 36.08 38.75
Kinyarwanda 40.84 43.87 45.04 47.56
Luganda 46.74 53.13 48.59 58.04
Luo 44.04 48.02 39.04 49.89
Mossi 26.05 26.42 30.96 31.71
Chichewa 53.72 52.93 48.52 51.94
Nigerian-Pidgin 66.71 67.50 39.96 81.91
Shona 43.47 43.87 34.37 46.58
Swahili 64.62 55.50 58.06 69.42
Setswana 51.45 54.05 39.83 56.97
Twi 35.62 45.69 28.54 51.10
Wolof 34.75 35.31 36.27 38.70
Xhosa 36.57 36.99 29.25 45.73
Yoruba 28.88 26.79 5.06 34.73
Zulu 38.07 34.61 28.34 36.50
[…]
Table 10: Results of cross-lingual evaluation on MasakhaNER 2.0.
Appendix D Zero-Shot Evaluation
Datasets
For the zero-shot evaluation, among NER benchmarks, we use $5$ English benchmarks from CrossNER Liu et al. (2021) and $2$ Chinese benchmarks, Weibo Peng and Dredze (2015) and Boson [^1] https://github.com/InsaneLife/. Among RE benchmarks, we adopt $2$ English benchmarks FewRel (Han et al., 2018), Wiki-ZSL (Chen and Li, 2021) and $3$ Chinese benchmarks, COAE2016222https://github.com/Sewens/COAE2016, IPRE (Wang et al., 2019), and SKE2020333https://aistudio.baidu.com/datasetdetail/177191. Among ED benchmarks, we use $3$ English benchmarks CrudeOil News Lee et al. (2021), RAMS Ebner et al. (2020), WikiEvents Li et al. (2021) and $2$ Chinese benchmarks FewFC (Zhou et al., 2021) and CCF law [^4] https://aistudio.baidu.com/projectdetail/4201483.
Results
Tables 11 and 12 show the zero-shot performance in English and Chinese, respectively. KnowCoder-X achieved SoTA performance across all benchmarks in Chinese tasks, demonstrating substantial and consistent improvements in generalization ability. In English tasks, KnowCoder-X exhibited a noteworthy average increase of $4.24$ points on the NER task. Similarly, in Chinese tasks, KnowCoder-X surpassed existing SoTA models on all benchmarks, achieving an average improvement of $3.29$ points. Moreover, the trends demonstrated in both Chinese and English show that the most significant improvement occurs in NER. We hypothesize that the NER alignment training in the first phase is more challenging to generalize to other IE tasks in out-of-domain scenarios, which leaves the introduction of RE and EE alignment training to future work.
Task Dataset GPT-4 YAYI-UIE IEPILE KnowCoder-X